# L’approximation arithmétique de Fujita sur courbes adéliques

### 27 Jan 2024 v1

@@ 2.1.3 Some Arakelov invariants
which is well-defined due to [7, §4.2]. This is independent of the choice of $s$ by the product formula of $S$.
We define
@@
We refer the readers to [7, §4.3] for more details and properties of these invariants.
@@ 2.2 Harder-Narasimhan filtration from $\mathbb{R}$-filtration
In this part, we introduce a method to filtrate an adelic vector bundle by its minimal slopes, which is called the Harder-Narasimhan filtration of an adelic vector bundle. See [3] for its origin and [7] for the version over adelic curve.
@@ 2.2.1 $\mathbb{R}$-filtrations on vector spaces
    $\mathcal{F}^{t}W=0$ for sufficiently large $t\in\mathbb{R}$, and $\mathcal{F}^{t}W=W$ for sufficiently small $t$;
@@ 2.2.2 A measure induced by $\mathbb{R}$-filtration
In fact, $\nu_{W}$ is a linear combination of some Dirac measure. More precisely, if the $\mathbb{R}$-filtration $\mathcal{F}$ consists of a flag
@@
$\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}}.$
In addition, we define the function
@@ 2.2.3 Harder-Narasimhan filtration
By the notion of $\mathbb{R}$-filtration, we introduce a filtration by minimal slopes of adelic bundles. This is original from an analogical study of vector bundles on a smooth projective curve [11].
@@
where $\Theta(E)$ denotes the set of all vector sub-spaces of $E$, and $\overline{F}$ is equipped with the sub-space norms. This $\mathbb{R}$-filtration $\mathcal{F}_{\operatorname{HN}}$ is called the Harder-Narasimhan filtration of $\overline{E}$.
The last equality in Definition 2.2.3 if from [7, Proposition 4.3.36]. In fact, the Harder-Narasimhan $\mathbb{R}$-filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
@@
In particular, still by [7, Proposition 4.3.36], we have $\mu_{n}=\widehat{\mu}_{\min}(\overline{E})$. The uniqueness of (3) is from [7, Theorem 4.3.58].
@@ 2.3.1 Approximable graded algebras
Let $K$ be a field, and $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be a graded $K$-algebra. We define $E_{\bullet}$ is approximable if:
@@
    $\liminf\limits_{n\to\infty}\frac{\dim\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim(E_{np})}>1-\epsilon,$
@@ 2.3.2 Convergence of measures on an approximable algebra
For a graded algebra, we equip an $\mathbb{R}$-filtration with all its homogeneous part. If it is an approximable graded algebras, we will have some properties of convergence.
@@
where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
@@
With all the above notations. Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
@@ 2.3.3 Limits of a measure from filtrated algebras
where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [2, §2.4] for more details.
@@ 3 Graded linear series of global section spaces
Let $S$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, and $\overline{L}$ be an adelic line bundle on $X$ over $S$. For our target, we will construct some graded linear series consisting of a particular kind of adelic vector bundles over $S$ from the above data.
@@ 3.1 Adelic line bundles on an arithmetic variety
In this part, we recall some notions about arithmetic positivity of the adelic line bundles on an arithmetic variety over adelic curve. We refer the readers to [9, §6 – §9] for a self-contained introduction to these notions.
@@ 3.1.1 Adelic line bundle
Let $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ be an adelic curve, and $X$ be a projective scheme over $\operatorname{Spec}K$. We say that $\overline{L}=(L,\varphi)$ is an adelic line bundle on $X$, if $L$ is a line bundle on $X$, and $\varphi=(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is a family of metrics on $L_{w}$, which is dominated and measurable.
In addition, we suppose that either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Then for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, we define
@@
as the value of $s$ at the point $x$ with respect to the metric, and
@@
We denote $\varphi=(\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}})_{w\in\Omega}$ the norm family on $H^{0}(X,L)$ over $S$ if there is no ambiguity. By [7, Theorem 6.1.13, Theorem 6.1.32], $(H^{0}(X,L),\varphi)$ is an adelic vector bundle over $S$.
Let $\overline{L}=(L,\varphi)$ and $\overline{M}=(M,\psi)$ be two adelic line bundles on $X$. We denote by $\overline{L}\otimes\overline{M}=(L\otimes M,\varphi+\psi)$ the adelic line bundle, where $\varphi+\psi$ is the tensor product of metrics. In particular, for $n\in\mathbb{N}$, we denote by $\overline{L}^{\otimes n}=(L^{\otimes n},n\varphi)$ the adelic line bundle of $n$-times tensor product of $\overline{L}$, and by $(H^{0}(X,L^{\otimes n}),n\varphi)$ the induced adelic vector bundle over $S$. By [7, Proposition 6.1.28, Proposition 6.1.31], all the above constructions are well-defined.
We refer the readers to [7, §6.1] for a precise explanation of §3.1.1, see also [9, §2.6 – 2.8].
@@ 3.1.2 Semi-positive metric
Let $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, $L$ be a line bundle on $X$, $\varphi=(\varphi_{w})_{w\in\Omega}$ and $\psi=(\psi_{w})_{w\in\Omega}$ be two families of metrics on $L$ such that $(L,\varphi)$ and $(L,\psi)$ are both adelic line bundles. Then we define the distance between $\varphi$ and $\psi$ as
@@
If $L$ is semiample, and if there exist a positive integer $m$ and a sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of quotient metric families (where $\varphi_{n}$ a metric family of $L^{\otimes mn}$), such that
@@ 3.1.3 Ampleness
In addition, an adelic line bundle $\overline{L}$ on $X$ is ample if it is relatively ample and if there exists an $\epsilon>0$ such that the inequality
@@
Compared with the arithmetically ample definition of S. Zhang in [26], an arithmetically ample Hermitian line bundle on an arithmetic variety requires the metrics on the line bundle to be smooth, while the definition in [9, Definition 9.1.1] over an adelic curve only require continuous metrics. In this article, we follow the definition given in [9, Definition 9.1.1].
@@ 3.1.4 Nefness
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say that $\overline{L}$ is relatively nef if there exists a relatively ample adelic line bundle $\overline{A}$ on $X$ and a positive integer $N$ such that, for any $n\geqslant N$, the tensor product $\overline{L}^{\otimes n}\otimes\overline{A}$ is relatively ample.
An adelic line bundle $\overline{L}$ is said to be nef or arithmetically nef if there exist an ample adelic line bundle $\overline{A}$ and a positive integer $M$ such that $\overline{L}^{\otimes m}\otimes\overline{A}$ is ample for any $m\geqslant M$ in the arithmetic sense.
@@ 3.1.5 Effectivity
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. If $s\in H^{0}(X,L)$ is a non-zero global section such that $\|s\|_{\varphi_{w}}\leqslant 1$ for any $w\in\Omega$, we say that the global section $s$ is effective. We say that $\overline{L}$ is effective or arithmetically effective if it admits at least one effective global section by the above definition.
@@ 3.2 Algebraic construction of graded linear series
In this part we only consider the geometric setting, which means everything works over an arbitrary field $K$. For every $D\in\mathbb{N}^{+}$, we consider the global section space $H^{0}(X,L^{\otimes D})$, and the graded $K$-algebra $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
@@ 3.2.1 Graded linear series containing an ample divisor
The following definition is original from [16, Definition 2.9], where it is called "Condition (C)". This notion plays a key role in the arithmetic Fujita approximation of [2].
@@
Let $X$ be a projective variety, and $L$ be a line bundle on $X$. Suppose that there exists a $p\in\mathbb{N}^{+}$, an ample line bundle $A$, an effective line bundle $M$ on $X$, and a non-zero section $s\in H^{0}(X,M)$, such that $L^{\otimes p}=A\otimes M$. We say a graded linear series $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ of $L$ contains an ample divisor if the homomorphism of graded algebras
@@
In §2.3, we have seen that an approximable graded linear series has many convergent properties. The following property is original from [16], which is applied in [2] to judge whether a graded linear series is approximable.
@@ 3.2.2 Big line bundles
If $W_{n}=H^{0}(X,L^{\otimes n})$, we denote $\operatorname{vol}(L)=\operatorname{vol}(W)$.
@@ 3.3 Arithmetic invariants of graded linear series
In this part, we consider arithmetic constructions.
@@
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity. By [7, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$. Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
Let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$. In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
@@
In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicify.
If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or we call it arithmetically big to distinguish it from the geometric case.
@@
For every $D\in\mathbb{N}$, Let $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E}_{D})\right)_{t\in\mathbb{R}}$ be the $\mathbb{R}$-filtration induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ following Definition 2.2.3, where we ignore the homogeneous degree $D\in\mathbb{N}$.
We denote by $\nu_{\overline{E}_{D}}$ the Borel probability measure induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ constructed in the sense of §2.2.2. Then we have the result below from Proposition 4.3.50, Proposition 4.3.51 and Corollary 4.3.52 of [7], see also [9, §2.5.4].

### 17 Mar 2024 v2

@@ 2.1.3 Some Arakelov invariants
which is well-defined due to [9, §4.2]. This is independent of the choice of $s$ by the product formula of $S$.
We define the positive Arakelov degree of $\overline{E}$ as
@@
We refer the readers to [9, §4.3] for more details and properties of these invariants.
@@ 2.2 Harder-Narasimhan filtration from $\mathbb{R}$-filtration
In this part, we introduce a method to filtrate an adelic vector bundle by its minimal slopes, which is called the Harder-Narasimhan filtration of an adelic vector bundle. We will use the language of $\mathbb{R}$-filtration to formulate this construction. See [5] for its origin and [9] for the version over adelic curve.
@@ 2.2.1 $\mathbb{R}$-filtrations on vector spaces
    $\mathcal{F}^{t}W=0$ for sufficiently large $t\in\mathbb{R}$, and $\mathcal{F}^{t}W=W$ for sufficiently small $t\in\mathbb{R}$;
@@ 2.2.2 A measure induced by $\mathbb{R}$-filtration
In fact, $\nu_{(W,\mathcal{F})}$ is a linear combination of some Dirac measures. More precisely, if the $\mathbb{R}$-filtration $\mathcal{F}$ consists of a flag
@@
$\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}},$
where $\delta_{\lambda_{i}}$ is the Dirac operator at the real number $\lambda_{i}$.
Next, we define the function
@@ 2.2.3 Harder-Narasimhan filtration
By the notion of $\mathbb{R}$-filtration, we introduce a filtration by the minimal slopes of adelic bundles. This is original from an analogical study of vector bundles on a smooth projective curve [13].
@@
where $\Theta(E)$ denotes the set of all vector subspaces of $E$, and $\overline{F}$ is equipped with the subspace norms. This $\mathbb{R}$-filtration $\mathcal{F}_{\operatorname{HN}}$ is called the Harder-Narasimhan filtration of $\overline{E}$.
The last equality in Definition 2.2.3 if from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
@@
In particular, still by [9, Proposition 4.3.36], we have $\mu_{n}=\widehat{\mu}_{\min}(\overline{E})$. The uniqueness of the flag (3) is from [9, Theorem 4.3.58].
@@ 2.3.1 Approximable graded algebras
Let $K$ be a field, and $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be a graded $K$-algebra. We define that $E_{\bullet}$ is approximable if:
@@
    $\liminf\limits_{n\to\infty}\frac{\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim_{K}(E_{np})}>1-\epsilon,$
@@ 2.3.2 Convergence of measures on an approximable algebra
For a graded algebra, we equip an $\mathbb{R}$-filtration with its each homogeneous part. If it is an approximable graded algebras, we will have some convergent properties.
@@
where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees of $x_{1}$ and $x_{2}$ are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
@@
With all the above notations. Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated approximable graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
@@ 2.3.3 Limits of a measure from filtrated algebras
where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [4, §2.4] for more details.
@@ 3 Graded linear series of global section spaces
Let $S$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, and $\overline{L}$ be an adelic line bundle on $X$ over $S$. For our purpose, we will construct some graded linear series consisting of some particular adelic vector bundles over $S$ from above data.
@@ 3.1 Adelic line bundles on an arithmetic variety
First, we recall some notions about arithmetic positivities of the adelic line bundles on arithmetic variety over an adelic curve, which are original from [8, 11]. We refer the readers to [11, §6 – §9] for a self-contained introduction to these notions.
@@ 3.1.1 Adelic line bundle
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $X$ be a projective scheme over $\operatorname{Spec}K$. We say that $\overline{L}=(L,\varphi)$ is an adelic line bundle on $X$, if $L$ is a line bundle on $X$, and $\varphi=(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is a family of metrics on $L_{w}$ for each $w\in\Omega$, which is dominated and measurable.
In addition, we suppose that either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Then we construct a family of norms on $H^{0}(X,L)$ in the way below: for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, define
@@
as the value of the global section $s$ at the point $x$ with respect to the given metric, and
@@
We denote $\varphi=(\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}})_{w\in\Omega}$ the norm family on $H^{0}(X,L)$ over $S$ if there is no ambiguity. By [9, Theorem 6.1.13, Theorem 6.1.32], $(H^{0}(X,L),\varphi)$ is an adelic vector bundle over $S$.
Let $\overline{L}=(L,\varphi)$ and $\overline{M}=(M,\psi)$ be two adelic line bundles on $X$. We denote by $\overline{L}\otimes\overline{M}=(L\otimes M,\varphi+\psi)$ the tensor adelic line bundle, where $\varphi+\psi$ is the tensor product of metrics. In particular, for $n\in\mathbb{N}$, we denote by $\overline{L}^{\otimes n}=(L^{\otimes n},n\varphi)$ the adelic line bundle of $n$-times tensor product of $\overline{L}$, and by $(H^{0}(X,L^{\otimes n}),n\varphi)$ the adelic vector bundle over $S$ defined above. By [9, Proposition 6.1.28, Proposition 6.1.31], all the above constructions are all well-defined.
We refer the readers to [9, §6.1] for a precise explanation of §3.1.1, see also [11, §2.6 – §2.8].
@@ 3.1.2 Semi-positive metric
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, $L$ be a line bundle on $X$, $\varphi=(\varphi_{w})_{w\in\Omega}$ and $\psi=(\psi_{w})_{w\in\Omega}$ be two families of metrics on $L$ such that $(L,\varphi)$ and $(L,\psi)$ are both adelic line bundles. Then we define the distance between $\varphi$ and $\psi$ as
@@
If $L$ is semi-ample, and if there exist a positive integer $m$ and a sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of quotient metric families (where $\varphi_{n}$ a metric family of $L^{\otimes mn}$), such that
@@ 3.1.3 Ampleness
In addition, an adelic line bundle $\overline{L}$ on $X$ is ample, if it is relatively ample and if there exists an $\epsilon>0$ such that the inequality
@@
We compare the above definition with the arithmetically ample definition of S. Zhang in [29], where an arithmetically ample Hermitian line bundle on an arithmetic variety requires the metrics on the line bundle to be smooth. The definition in [11, Definition 9.1.1] over an adelic curve only require continuous metrics. In this article, we follow the definition given in [11, Definition 9.1.1] as above.
@@ 3.1.4 Nefness
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say that $\overline{L}$ is relatively nef if there exists a relatively ample adelic line bundle $\overline{A}$ on $X$ and a positive integer $N$ such that, for any $n\geqslant N$, the tensor product of adelic line bundles $\overline{L}^{\otimes n}\otimes\overline{A}$ is relatively ample.
An adelic line bundle $\overline{L}$ is said to be nef or arithmetically nef, if there exist an arithmetically ample line bundle $\overline{A}$ and a positive integer $M$ such that $\overline{L}^{\otimes m}\otimes\overline{A}$ is ample for any $m\geqslant M$ in the arithmetic sense.
@@ 3.1.5 Effectivity
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is effective.
We say that an adelic line bundle $\overline{L}$ is effective or arithmetically effective, if it admits at least one effective global section by the above definition.
@@ 3.2 Algebraic construction of graded linear series
In this part we will only consider the geometric setting, which means everything works over an arbitrary field with no metric. For every $D\in\mathbb{N}^{+}$, we consider the global section space $H^{0}(X,L^{\otimes D})$, and the graded $K$-algebra $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
@@ 3.2.1 Graded linear series containing an ample divisor
The following definition is original from [18, Definition 2.9], where it is called "Condition (C)". This notion plays a key role in the arithmetic Fujita approximation of [4], and we follow the notation of [4, Definition 3.1].
@@
Let $X$ be a projective variety, and $L$ be a line bundle on $X$. Suppose that there exists a $p\in\mathbb{N}^{+}$, an ample line bundle $A$, an effective line bundle $M$ on $X$, and a non-zero section $s\in H^{0}(X,M)$, such that $L^{\otimes p}=A\otimes M$. In this case, we say a graded linear series $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ of $L$ contains an ample divisor if the homomorphism of graded algebras
@@
In §2.3.2, we have seen that a filtrated approximable graded linear series has many convergent properties. The following property is original from [18], which is applied in [4] to judge whether a graded linear series is approximable.
@@ 3.2.2 Big line bundles
If $W_{n}=H^{0}(X,L^{\otimes n})$, we denote $\operatorname{vol}(L)=\operatorname{vol}(W_{\bullet})$.
@@ 3.3 Arithmetic invariants of graded linear series
In this part, we consider some arithmetic constructions.
@@
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity. By [9, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$. Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$. In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
@@
In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.
If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or arithmetically big to distinguish it from the geometric case.
@@
For every $D\in\mathbb{N}$, let $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E}_{D})\right)_{t\in\mathbb{R}}$ be the $\mathbb{R}$-filtration induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ following Definition 2.2.3, where we ignore the homogeneous degree $D\in\mathbb{N}$ of $\overline{E}_{D}$.
We denote by $\nu_{\overline{E}_{D}}$ the Borel probability measure induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ constructed in the sense of §2.2.2. Then we have the result below from Proposition 4.3.50, Proposition 4.3.51 and Corollary 4.3.52 of [9], see also [11, §2.5.4].